Database last updated on Thu, 05 Feb 2026 20:10:52 +0000 by localhost

Critical warnings from b-files' comments for days between 2022084 and 2022084 (0 sec, 9 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2022 084 20 32 43 Motor #3: appears stuck with a reference sensor activated.21B08422.191TORON#19105:26:151-1
rl:2022 084 20 34 52 New value set for watermark SUPPLY.VOLTAGE.HIGH[-5VSEC]: -4.333701 Volts.123B08422.191TORON#19105:26:231-1
rl:2022 085 00 09 23 PMT counter failure11B08422.191TORON#19105:27:301-1
td:Date changed from 24/03/22 to 25/03/2211B08422.191TORON#19115:15:241-1
hp:Spectrometer alignment trouble.41B08422.191TORON#19115:23:281-1
hg:Brewer failed to respond 5 times51B08422.191TORON#19119:07:551-1
hg:Lamp not on... hg test terminated F1= 0 F4= 021B08422.191TORON#19120:26:101-1
hg:Lamp not detected f1= 0 f4= 021B08422.191TORON#19120:26:111-1
hg:No intensity detected. Check PMT.21B08422.191TORON#19120:26:111-1

Instrument constant changes for the week before 2022084 (3 sec, 0 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
191202208465.6

Non-critical warnings from b-files' comments for days between 2022084 and 2022084 (0 sec, 11 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2022 084 20 32 42 Reset No -> Tepid from 2022 084 20 26 30 until 2031 048 14 04 21 . Counts: 1,013B08422.191TORON#19105:26:101-1
rl:2022 084 20 32 43 Checksum bad in low level command.13B08422.191TORON#19105:26:121-1
rl:2022 084 20 32 43 Motor #3: appears stuck with a reference sensor activated.13B08422.191TORON#19105:26:151-1
rl:2022 084 20 32 43 Motor #11: appears stuck with a reference sensor activated.13B08422.191TORON#19105:26:181-1
rl:2022 084 20 34 52 MAIN.C : motor initialization timed out.25B08422.191TORON#19105:26:201-1
rl:2022 084 21 00 47 Reset Warm -> Warm from 2022 084 21 00 43 until 2022 084 21 00 44 . Counts: 0,143B08422.191TORON#19105:26:581-1
rl:2022 084 21 04 04 Bad command string. Command string buffer cleared.43B08422.191TORON#19105:27:031-1
rl:2022 084 21 04 04 Bad command string. Command string buffer cleared.43B08422.191TORON#19105:27:031-1
rl:2022 084 21 04 04 <=43B08422.191TORON#19105:27:041-1
hp:HOME key pressed299B08422.191TORON#19120:13:481-1
hp:Going to menu299B08422.191TORON#19120:14:201-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhghprssl
084TORONTO#191inside,Must do HG now,Waiting until ZA = 170 now 126.5,Waiting until ZA = 170 now 127.6,Waiting until ZA = 170 now 128.6,Waiting until ZA = 170 now 129.6,Waiting until ZA = 170 now 130.5,Waiting until ZA = 170 now 131.3,Waiting until ZA = 170 now 132,Waiting until ZA = 170 now 132.6,Waiting until ZA = 170 now 133,Waiting until ZA = 170 now 133.4,Waiting until ZA = 170 now 133.7,Waiting until ZA = 170 now 133.9,Waiting until ZA = 170 now 1343213735
(1 sec)

Relative and absolute humidity for day 2022084 (1 sec, 1 rows)

SerialRH, %AH, g/m37-day RH historyDes-nt replcd days ago
1917.61.9